# Optimizing the pump coupling for a three-wave mixing Josephson parametric amplifier

### v1

@@ Abstract
Josephson element-based parametric amplifiers (JPAs) typically require rf pump power that is several orders of magnitude stronger than the maximum signal power they can handle. The low power efficiency and strong pump leakage towards signal circuitry could be critical concerns in application. In this work, we discuss how to optimize the pump coupling scheme for a three-wave mixing JPA by employing microwave filtering techniques, with the goal of maximizing the pump power efficiency and minimize pump leakage without sacrificing other properties of interest. We implement the corresponding filter design in a SNAIL-based JPA and demonstrate more than three orders of magnitude improvement in both power efficiency and pump leakage suppression compared to a similar device with regular capacitive coupling, while maintaining state-of-the-art dynamic range and near-quantum-limited noise performance. Furthermore, we show experimentally that the filter-coupled JPA is more robust against noise input from the pump port, exhibiting no significant change in added noise performance with up to 4 K of effective noise temperature at the pump port.
@@ I Introduction
Josephson element-based parametric amplifiers (JPAs)[1] have become critical tools in quantum measurements in which the information is carried by a few microwave photons. These amplifiers provide large gain with near quantum-limited added noise, enabling highly efficient microwave measurements in single-shot qubit readout[2], electron spin resonance detection[3], and axion search[4]. Recent research has focused on improving JPA bandwidth and compression power[5, 6, 7, 8, 9, 10, 11, 12], which is necessary for multiplexed qubit readout[13, 14] in large-scale quantum processors. However, a major challenge in operating a large array of JPAs at millikelvin temperature is the heat load added to the cryostat by the microwave drives applied for amplification, commonly referred to as the ‘pump’.
@@
To achieve quantum-limited noise performance with a JPA, the pump is typically delivered to the JPA through a heavily attenuated transimission line to reduce thermal noise. In addition, the pump line is weakly coupled to the JPA, usually with a small capacitance or mutual inductance, to minimize the signal photon loss. As a result of these choices, the pump power required by the JPA is usually several orders of magnitude higher than the signal power. For instance, a state-of-the-art 3-wave-mixing JPA dissipates around 0.01 to 1 µW of power on the base plate of a dilution refrigerator, whose cooling power is typically 10 to 100 µW at 20 mK [15]. This is a practical limiting factor in scaling up superconducting quantum processors.
@@
In addition, the strong pump power incident on the JPA also poses challenges for protecting qubits against unwanted back-action. Since the coupling strength on the signal port is much higher than that on the pump port, the intra-amplifier pump photons predominantly exit the amplifier through the signal port and propagate towards the signal source. It usually requires multiple non-reciprocal components (other than the circulator shown in Fig. 1) or filters to achieve sufficient isolation of the pump leakage from the signal source. These elements reduce readout efficiency due to their insertion loss, and can emit thermal photons that dephase the qubit due to heating by the pump leakage.
@@
Figure 1: Minimal schematic describing a reflection JPA as the first stage of the amplifier chain, illustrating the problems addressed in this article. Red arrows indicate information-carrying signal and blue arrows indicate pump. There is generally a trade-off between signal loss and the required pump power at port 2. Furthermore, one circulator at port 1 (for signal directionality) is not sufficient to isolate pump leakage from the input circuitry.
@@
Therefore, reducing the pump power needed to operate JPAs is crucial for deploying large-scale systems and for improving the readout performance and qubit coherence. In addition, broadband and higher compression power JPAs for multiplexed qubit readout often demands elevated intra-amplifier pump power [10, 16, 15, 17, 12]. Hence, it is desirable to increase pump power coupling at the pump port while further suppressing pump leakage from the signal port.
@@
JPAs with pump tones far detuned from the signal frequency offer the opportunity to address these issues by applying microwave filtering techniques. In this work, we demonstrate near 30 dB reduction in both pump power and pump leakage in a three-wave mixing SNAIL parametric amplifier (SPA) by upgrading its capacitively coupled ports with on-chip filters, while retaining its near quantum-limited noise performance. These improvements would enable deployment of a large bank of parallel SPAs with minimal heat load added to the cryostat, facilitating scaling up of superconducting quantum processors. The technique introduced in this work is also applicable to 4-wave-mixing parametric amplifiers with sideband pumping[18], and more generally to other parametric processes activated by off-resonant pumps.
@@
This article is organized as followed: in Section II we introduce power efficiency as the metric for characterizing the efficiency of converting pump power to output signal power of a JPA and formulate its optimization as a specific circuit synthesis problem. In Section III we demonstrate a realization of a filter-coupled SPA that fulfills the optimization goals. In Section IV we compare the filter-coupled SPA with a standard capacitor-coupled SPA and demonstrate improvements in both power efficiency and pump leakage suppression. In Section V, we show the improved noise rejection at signal frequency on the pump port of the filter-coupled SPA, which would allow the amplifier to achieve near quantum-limited noise performance with reduced attenuation on the pump line and further decreases its head load to the cryostat.
@@ II Optimization of JPA power efficiency
A parametric amplifier converts power from its pump into the output signal. The figure-of-merit that characterizes the efficiency of this process is the power efficiency, $\eta_{p}$ defined as the ratio between the output signal power at 1-dB gain compression point ($P^{\mathrm{out}}_{1\mathrm{dB}}$), and the corresponding pump power ($P_{\mathrm{pump}}$):
$\eta_{p}=\frac{P^{\mathrm{out}}_{1\mathrm{dB}}}{P_{\mathrm{pump}}}$ (1)
JPAs typically have power efficiency ranging from $10^{-2}$ to $10^{-6}$, as reviewed in [15]. The power efficiency of 3-wave-mixing JPAs [19, 20, 21, 22, 7, 8, 17], due to the large frequency detuning of the pump, is generally orders of magnitude lower than that of their 4-wave-mixing counterparts [10, 5, 22]. In this section, we formulate the goals and constraints for optimizing JPA power efficiency within the framework of network synthesis.
We limit our discussion to the current-pumped JPA whose nonlinearity originates from a lumped Josephson dipole, namely a two-terminal Josephson element. In particular, we consider a 3-wave-mixing JPA operating in the phase-preserving mode, in which the wave components at pump frequency $\omega_{p}$, signal frequency $\omega_{s}$, and idler frequency $\omega_{i}=\omega_{p}-\omega_{s}$ are coupled via the 3rd-order nonlinearity of the Josephson dipole.
As illustrated in Fig. 2(a), we view the linear embedding circuit of the Josephson dipole in the JPA as a microwave network with two external ports connected to the signal and pump sources respectively, and one internal port loaded by the Josephson dipole. We perform signal-flow analysis on the system at pump and signal frequencies separately, in order to formulate the power efficiency and amplification performance in terms of the circuit parameters.
Figure 2: Modeling the JPA as a scattering system. (a) The JPA consists of a lumped Josephson dipole embedded in a linear network, which has two external ports and one internal port. The external ports - port 1 and port 2 - connect to the signal and pump source respectively. The internal port - port 3 - is loaded by the Josephson dipole. (b) The pump at frequency $\omega_{p}$ is applied from port 2, creating a voltage phasor $V_{p}^{\mathrm{J}}$ at the Josephson dipole. (c) At signal frequency $\omega_{s}$, the Josephson dipole produces an additive admittance arising from pumping. Reflection gain for the signal is measured from port 1.
To relate the power efficiency to the circuit parameters in Fig. 2(b), we first rewrite the definition (1) as
@@
where $P^{\mathrm{J}}_{p}=\left|V^{\mathrm{J}}_{p}\right|^{2}/\omega_{p}L_{J}$ is the steady-state pump power on the Josephson dipole created by the pump tone applied on port 2. The first factor in Eq. (2) is the ratio between the output signal power at 1 dB gain compression point and the pump power incident onto the Josephson dipole. We will refer to this factor as the nonlinear mixing efficiency $\eta^{\mathrm{nl}}_{p}$. The second factor in Eq. (2) is the ratio between the pump power incident onto the Josephson dipole and the power $P_{\mathrm{pump}}=|V^{\mathrm{in},2}_{p}|^{2}/Z_{0}$ incident on the device pump port, which we will refer to as the pump coupling efficiency $\eta^{\mathrm{c}}_{p}$. Hence $\eta_{p}=\eta^{\mathrm{nl}}_{p}\eta^{\mathrm{c}}_{p}$.
The power efficiency of a JPA can be improved by increasing the nonlinear mixing efficiency $\eta^{\mathrm{nl}}_{p}$ and/or the pump coupling efficiency $\eta^{\mathrm{c}}_{p}$. Improving $\eta_{p}$ by increasing $\eta^{\mathrm{nl}}_{p}$ has been demonstrated in experiment. For example, by operating a 3-wave-mixing JPA at its ‘Kerr-free’ point, where the 4th-order nonlinearity is minimized, the $P^{\mathrm{out}}_{1\mathrm{dB}}$ can be improved by more than 10 dB under similar pump power [15]. Recently, a general method for engineering higher-order nonlinearities of Josephson elements for better JPA power efficiency has been proposed [23]. However, it is important to note that getting $\eta^{\mathrm{nl}}_{p}$ close to unity leads to the effect of pump depletion[1, 24], causing the device to deviate from the linear amplifier regime. More investigation is required for a JPA operated in the pump depletion regime.
In this work, we instead focus on improving $\eta_{p}$ by maximizing $\eta^{\mathrm{c}}_{p}$, which amounts to establishing the desired $P^{\mathrm{J}}_{p}$ across the Josephson dipole with minimal pump power $P_{\mathrm{pump}}$. In contrast to $\eta^{\mathrm{nl}}_{p}$, engineering $\eta^{\mathrm{c}}_{p}$ to its theoretical upper limit of unity does not lead to pump depletion.
@@
in which $P_{\mathrm{leak}}=|V^{\mathrm{out},1}_{p}|^{2}/Z_{0}$ is the pump leakage power. Therefore, the pump leakage power scales as and is upper bounded by $P^{\mathrm{J}}_{p}/\eta_{p}^{\mathrm{c}}$, and improving $\eta_{p}^{\mathrm{c}}$ will generally reduce pump leakage. We will discuss further suppression of pump leakage in Section III and Section IV by filtering on the signal port.
The pump coupling efficiency of the JPA is obtained by analyzing the circuit in Fig. 2(b), where we approximate the nonlinear Josephson dipole as a linear inductance $L_{J}$. The voltage across the Josephson dipole created by the pump, $V^{\mathrm{J}}_{p}$, can be evaluated with respect to the input phasor $V^{\mathrm{in},2}_{p}$. The pump coupling efficiency is thus given by:
$\eta_{p}^{\mathrm{c}}=\left|\frac{1+\Gamma_{3}[\omega_{p}]}{1-\Gamma_{3}[\omega_{p}]S_{33}[\omega_{p}]}S_{32}[\omega_{p}]\right|^{2}\frac{Z_{0}}{\omega_{p}L_{\mathrm{J}}}$ (4)
where $\Gamma_{3}[\omega_{p}]=\frac{j\omega_{p}L_{J}-Z_{0}}{j\omega_{p}L_{J}+Z_{0}}$ is the reflection coefficient from the linear inductance $L_{J}$ loading port 3 of the network, $S_{33}[\omega_{p}]$ and $S_{32}[\omega_{p}]$ are the scattering parameters of the network. Therefore, maximizing the pump coupling efficiency amounts to a linear network synthesis problem.
@@
We now discuss the requirements for the circuit to operate as a resonant amplifier. First, a standing wave eigenmode must exist in the circuit. This requirement is captured by the following Laplace domain equation:
$Z^{\mathrm{Th}}(s)+sL_{J}=0$ (5)
@@
where $Z^{\mathrm{Th}}=Z_{0}(1+S_{33})/(1-S_{33})$ is the Thevenin impedance viewed by the Josephson dipole. The solution of Eq. (5), $s_{0}=j\omega_{0}-\frac{\kappa_{0}}{2}$, provides the resonance frequency $\omega_{0}$ and external coupling rate $\kappa_{0}$ for the electromagnetic eigenmode of the system. Eq. (5) and its reformulation in the Fourier domain are discussed in Appendix C.
Second, signal power gain is expected in reflection from port 1, which is given by:
@@
where $\Gamma_{3}[\omega_{s}]$ is the reflection coefficient loading port 3 at the signal frequency controlled by the pump. Under the stiff pump approximation[1], $\Gamma_{3}[\omega_{s}]$ from a current-pumped Josephson dipole is obtained from a linearized model similar to the flux-pumped case [25, 26, 27] (see Appendix D for details). In our model, as illustrated in Fig. 2(c), the response of the Josephson dipole at signal frequency is represented by the inductance $L_{J}$ in parallel with an effective admittance $Y^{\mathrm{nl}}(V^{\mathrm{J}}_{p})$ arising from pumping. This gives rise to the reflection coefficient:
$\Gamma_{3}[\omega_{s}]=\frac{1-Z_{0}\left(\frac{1}{j\omega_{s}L_{J}}+Y^{\mathrm{nl}}(V^{\mathrm{J}}_{p})\right)}{1+Z_{0}\left(\frac{1}{j\omega_{s}L_{J}}+Y^{\mathrm{nl}}(V^{\mathrm{J}}_{p})\right)}$ (7)
Amplification for the signal requires $\left|\Gamma_{3}[\omega_{s}]\right|>1$, which translates into $Y^{\mathrm{nl}}(V^{\mathrm{J}}_{p})$ having a negative real part. Moreover, achieving high gain ($G\gg 1$) requires $1-\Gamma_{3}[\omega_{s}]S_{33}[\omega_{s}]\approx 0$, which is satisfied for near-resonance signals ($\omega_{s}\approx\omega_{0}$), as explained in Appendix D.
Lastly, to achieve a noise performance approaching the quantum limit, it requires minimizing the signal loss via port 2. The transmission coefficient from port 1 to port 2 at signal frequency is given by:
$T=\left|\frac{V^{\mathrm{out},2}_{s}}{V^{\mathrm{in},1}_{s}}\right|^{2}=\left|S_{21}[\omega_{s}]+\frac{S_{23}[\omega_{s}]\Gamma_{3}[\omega_{s}]S_{31}[\omega_{s}]}{1-\Gamma_{3}[\omega_{s}]S_{33}[\omega_{s}]}\right|^{2}$ (8)
@@
Near-quantum-limited noise performance for reflection gain requires $T\ll G$, essentially demanding $\left|S_{23}[\omega_{s}]\right|\ll\left|S_{13}[\omega_{s}]\right|$. In the standard design with capacitively coupled ports, for example, this condition is satisfied by having a large coupling capacitance at port 1 and a small coupling capacitance at port 2. This choice, however, leads to poor transmission at pump frequency from port 2 as well, conflicting with the demand for large pump coupling efficiency given in Eq. (4).
@@
The frequency separation between the pump and signal for the 3-wave-mixing process provides an opportunity to satisfy these seemingly conflicting requirements. In the following sections, we demonstrate that the optimization goals at pump frequency $\omega_{p}$ and signal frequency $\omega_{s}$ can be achieved simultaneously by engineering the coupling network to be strongly frequency selective on a 3-wave-mixing JPA.

### v2

@@ Abstract
Josephson parametric amplifiers (JPAs) typically require rf pump power that is several orders of magnitude stronger than the maximum signal power they can handle. The low power efficiency and strong pump leakage towards signal circuitry could be critical concerns in application. In this work, we discuss how to optimize the pump coupling scheme for a three-wave mixing JPA by employing microwave filtering techniques, with the goal of maximizing the pump power efficiency and minimize pump leakage without sacrificing other properties of interest. We implement the corresponding filter design in a SNAIL-based JPA and demonstrate more than three orders of magnitude improvement in both power efficiency and pump leakage suppression compared to a similar device with conventional capacitive coupling, while maintaining state-of-the-art dynamic range and near-quantum-limited noise performance. Furthermore, we show experimentally that the filter-coupled JPA is more robust against noise input from the pump port, exhibiting no significant change in added noise performance with up to 4 K of effective noise temperature at the pump port.
@@ I Introduction
Josephson parametric amplifiers (JPAs) Roy and Devoret (2018) have become critical tools in quantum measurements in which the information is carried by a few microwave photons. These amplifiers provide large gain with near-quantum-limited added noise, enabling highly efficient microwave measurements in single-shot qubit readout Jeffrey et al. (2014), electron spin resonance detection Bienfait et al. (2016), and axion search Brubaker et al. (2017). Recent research has focused on improving JPA bandwidth and compression power Roy et al. (2015); Eichler and Wallraff (2014); Liu et al. (2017); Frattini et al. (2018); Naaman et al. (2019); Planat et al. (2019); Kaufman et al. (2023, 2024), which is necessary for multiplexed qubit readout Kundu et al. (2019); White et al. (2023) in large-scale quantum processors; however, one operational challenge in scaling up superconducting quantum processors is the heat load added to the cryostat by the microwave drives applied to the JPAs, commonly referred to as the ‘pump’.
@@
To achieve quantum-limited noise performance with a JPA, the pump is typically delivered to the JPA through a heavily attenuated transmission line to reduce thermal noise. In addition, the pump line is weakly coupled to the JPA, usually with a small capacitance or mutual inductance, to minimize the signal photon loss. As a result of these choices, the pump power required by the JPA is usually several orders of magnitude higher than the signal power. For instance, a state-of-the-art three-wave-mixing JPA dissipates around 0.01 to 1 µW of power on the base plate of a dilution refrigerator, whose cooling power is typically 10 to 100 µW at 20 mK Sivak et al. (2019). This is a practical limiting factor in operating a large array of JPAs at millikelvin temperatures.
@@
In addition, the strong pump power incident on the JPA also poses challenges for protecting qubits against unwanted back-action Kaufman et al. (2024). Since the coupling strength on the signal port is much higher than that on the pump port, the intra-amplifier pump photons predominantly exit the amplifier through the signal port and propagate toward the signal source. Multiple nonreciprocal components (other than the circulator shown in Fig. 1) or filters are usually required to achieve sufficient isolation of the pump leakage from the signal source. These elements reduce readout efficiency due to their insertion loss, and they can emit thermal photons that dephase the qubit due to heating by the pump leakage.
@@
Figure 1: Minimal schematic describing a reflection JPA as the first stage of the amplifier chain, illustrating the problems addressed in this article. Red arrows indicate information-carrying signals and blue arrows indicate pump waveforms. There is generally a trade-off between signal loss and the required pump power at port 2. Furthermore, one circulator at port 1 (for signal directionality) is not sufficient to isolate pump leakage from the input circuitry.
@@
Reducing the pump power needed to operate JPAs is therefore crucial for two reasons: reducing the heat load and improving both the readout performance and qubit coherence. These issues become increasingly pressing because broadband, higher-saturation-power JPAs often require elevated intra-amplifier pump power  Planat et al. (2019); Liu et al. (2020); Sivak et al. (2020); Kaufman et al. (2024) to operate. This therefore points to a clear design objective to increase pump-power coupling at the pump port while further suppressing pump leakage at the signal port.
@@
JPAs with pump tones far detuned from the signal frequency offer the opportunity to address these issues by applying microwave filtering techniques. In this work, we demonstrate a near-30-dB reduction in both pump power and pump leakage in a three-wave-mixing SNAIL parametric amplifier (SPA) by upgrading its capacitively coupled ports with on-chip filters, while retaining its near-quantum-limited noise performance. These improvements would enable deployment of a large bank of parallel SPAs with minimal heat load added to the cryostat, facilitating the scaling up of superconducting quantum processors. The technique introduced in this work is also applicable to four-wave-mixing parametric amplifiers with sideband pumping Kamal et al. (2009), and more generally to other parametric processes activated by off-resonant pumps.
@@
The remainder of this article is organized as follows: In Section II, we introduce power efficiency as the metric for characterizing the efficiency of converting pump power to the output signal power of a JPA, and we formulate its optimization as a specific circuit synthesis problem. In Section III, we demonstrate a realization of a filter-coupled SPA that fulfills the optimization goals. In Section IV, we compare the filter-coupled SPA with a control device with conventional capacitive coupling, demonstrating improvements in both power efficiency and pump-leakage suppression. In Section V, we show the improved noise rejection at the signal frequency on the pump port of the filter-coupled SPA, which would allow the amplifier to achieve near-quantum-limited noise performance with reduced attenuation on the pump line and further decrease its heat load to the cryostat.
@@ II Optimization of JPA power efficiency
A parametric amplifier converts power from its pump into an output signal. The figure of merit that characterizes the efficiency of this process is the power efficiency $\eta_{p}$, which is defined as the ratio between the output signal power at 1-dB gain compression point ($P^{\mathrm{out}}_{1\mathrm{dB}}$) and the corresponding pump power ($P_{\mathrm{pump}}$):
$\eta_{p}=\frac{P^{\mathrm{out}}_{1\mathrm{dB}}}{P_{\mathrm{pump}}}.$ (1)
JPAs typically have power efficiencies ranging from $10^{-2}$ to $10^{-6}$. As reviewed in Ref. Sivak et al. (2019), the power efficiency of three-wave-mixing JPAs is generally orders of magnitude lower than that of their four-wave-mixing counterparts. In the current-pumping scheme Liu et al. (2017); Frattini et al. (2018); Planat et al. (2019); Sivak et al. (2020), the pump of a three-wave-mixing JPA is handicapped in coupling strength due to a large frequency detuning from resonance. In contrast, in the flux-pumping scheme, the pump coupling is insensitive to detuning since a differential mode is driven, which is also advantageous in terms of suppressing pump leakage Yamamoto et al. (2008); However, it is challenging to engineer a strong mutual-inductive coupling. As a result, the reported power efficiencies for flux-pumped JPAs are typically at the $10^{-5}$ level Simoen et al. (2015); Zhou et al. (2014); Elo et al. (2019); Mutus et al. (2013), although there is ongoing research on improving these values  Urade et al. (2021). In this section, we formulate the goals and constraints for optimizing the power efficiency of current-pumped JPAs within a network synthesis framework.
We consider a three-wave-mixing JPA operating in the phase-preserving mode, in which the wave components at pump frequency $\omega_{p}$, signal frequency $\omega_{s}$, and idler frequency $\omega_{i}=\omega_{p}-\omega_{s}$ are coupled via the third-order nonlinearity. We focus on a current-pumped JPA whose nonlinearity originates from a lumped Josephson dipole, namely a two-terminal Josephson element. As illustrated in Fig. 2(a), we view the linear embedding circuit of the Josephson dipole in the JPA as a microwave network with two external ports connected to the signal and pump sources respectively, and one internal port loaded by the Josephson dipole. Separate signal-flow analyses are performed at the pump and signal frequencies to formulate the power efficiency and amplification performance in terms of the circuit parameters.
Figure 2: Modeling the JPA as a scattering system. (a) The JPA consists of a lumped Josephson dipole embedded in a linear network, which has two external ports and one internal port. The external ports — port 1 and port 2 — connect to the signal and pump source, respectively. The internal port — port 3 — is loaded by the Josephson dipole. (b) The pump at frequency $\omega_{p}$ is applied from port 2, creating a voltage phasor $V_{p}^{\mathrm{J}}$ across the Josephson dipole. (c) At signal frequency $\omega_{s}$, the Josephson dipole produces an additive admittance arising from pumping. The reflection gain for the signal is measured from port 1.
We decompose the power efficiency into two meaningful factors by rewriting the definition in Eq. (1) as
@@
where $P^{\mathrm{J}}_{p}=\left|V^{\mathrm{J}}_{p}\right|^{2}/\omega_{p}L_{\mathrm{J}}$ is the steady-state pump power resulting from the rms voltage $\left|V_{P}^{\mathrm{J}}\right|$ across the Josephson dipole. The first factor in Eq. (2) is the ratio between the output signal power at 1-dB gain compression point and the pump power incident on the Josephson dipole. We will refer to this factor as the nonlinear mixing efficiency $\eta^{\mathrm{nl}}_{p}$. The second factor in Eq. (2) is the ratio between the pump power incident on the Josephson dipole and the power $P_{\mathrm{pump}}=|V^{\mathrm{in},2}_{p}|^{2}/Z_{0}$ incident on the device pump port (port 2), which we will refer to as the pump coupling efficiency $\eta^{\mathrm{c}}_{p}$. Hence $\eta_{p}=\eta^{\mathrm{nl}}_{p}\eta^{\mathrm{c}}_{p}$. Under the stiff-pump approximation Roy and Devoret (2018), these two factors are independent of each other.
The power efficiency of a JPA can be improved by increasing the nonlinear mixing efficiency $\eta^{\mathrm{nl}}_{p}$ and/or the pump coupling efficiency $\eta^{\mathrm{c}}_{p}$. Improving $\eta_{p}$ by increasing $\eta^{\mathrm{nl}}_{p}$ has been demonstrated in experiments. For example, by operating a three-wave-mixing JPA at its ‘Kerr-free’ point, where the fourth-order nonlinearity is minimized, the $P^{\mathrm{out}}_{1\mathrm{dB}}$ value can be improved by more than 10 dB under similar pump power  Sivak et al. (2019). Recently, a general method for engineering higher-order nonlinearities of Josephson elements for better JPA power efficiency has been proposed  Hougland et al. (2025). However, it is important to note that obtaining a $\eta^{\mathrm{nl}}_{p}$ value close to unity leads to the effect of pump depletion Roy and Devoret (2018); Schackert et al. (2013), causing the device to deviate from the linear amplifier regime. More investigation is required for a JPA operated in the pump-depletion regime.
In this work, we focus on improving $\eta_{p}$ by increasing the pump coupling efficiency. The goal is to establish the desired $P^{\mathrm{J}}_{p}$ across the Josephson dipole with minimal pump power $P_{\mathrm{pump}}$. This goal amounts to maximizing:
$\eta_{p}^{\mathrm{c}}=\left|\frac{V_{p}^{\mathrm{J}}}{V^{\rm{in},2}_{p}}\right|^{2}\frac{Z_{0}}{\omega_{p}L_{\mathrm{J}}}$ (3)
for the circuit shown in Fig. 2(b), where the nonlinear Josephson dipole is approximated as a linear inductance $L_{\mathrm{J}}$ at the pump frequency. Optimizing the pump coupling is thus formulated as a linear network synthesis problem. Notably, in contrast to the nonlinear mixing efficiency $\eta^{\mathrm{nl}}_{p}$, engineering $\eta^{\mathrm{c}}_{p}$ to its theoretical upper limit of unity does not lead to pump depletion.
@@
in which $P_{\mathrm{leak}}=|V^{\mathrm{out},1}_{p}|^{2}/Z_{0}$ is the pump leakage power. Therefore, the pump leakage power scales as and is upper bounded by $P^{\mathrm{J}}_{p}/\eta_{p}^{\mathrm{c}}$, and improving $\eta_{p}^{\mathrm{c}}$ will generally reduce the pump leakage. We will discuss further suppression of pump leakage in Secs. III and IV by filtering at the signal port.
@@
The maximization of Eq. (3) is subject to constraints required for the circuit to function as a resonant amplifier. First, a standing wave eigenmode must exist in the circuit. This requirement is captured by the Laplace-domain equation
$Z^{\mathrm{Th}}(s)+sL_{\mathrm{J}}=0$ (5)
@@
where $Z^{\mathrm{Th}}=Z_{0}(1+S_{33})/(1-S_{33})$ is the Thevenin impedance viewed by the Josephson dipole. The solution to Eq. (5), $s_{0}=j\omega_{0}-\frac{\kappa_{0}}{2}$, provides the resonance frequency $\omega_{0}$ and external coupling rate $\kappa_{0}$ for the electromagnetic eigenmode of the system. Eq. (5) and its reformulation in the Fourier domain are discussed in Appendix C.
@@
Second, signal amplification requires the reflection coefficient from the Josephson dipole to have a magnitude greater than unity. Under the stiff-pump approximation, the response of the Josephson dipole at the signal frequency can be modeled as the inductance $L_{\mathrm{J}}$ in parallel with an effective admittance $Y^{\mathrm{nl}}(V^{\mathrm{J}}{p})$ arising from pumping, as illustrated in Fig. 2(c). This effective admittance representation resembles the ‘pumpistor’ model developed for flux-pumped Josephson parametric devices Sundqvist et al. (2013); Mutus et al. (2013); Sundqvist and Delsing (2014); Naaman and Aumentado (2022), with details provided in Appendix D. The reflection coefficient at the signal frequency is therefore given by:
$\Gamma_{3}[\omega_{s}]=\frac{1-Z_{0}(\frac{1}{j\omega_{s}L_{\mathrm{J}}}+Y^{\mathrm{nl}}(V^{\mathrm{J}}{p}))}{1+Z_{0}(\frac{1}{j\omega_{s}L_{\mathrm{J}}}+Y^{\mathrm{nl}}(V^{\mathrm{J}}{p}))}$ (6)
The requirement for amplification $\left|\Gamma_{3}[\omega_{s}]\right|>1$ is satisfied when $Y^{\mathrm{nl}}(V^{\mathrm{J}}{p})$ exhibits a negative real part.
@@
Lastly, for near-quantum-limited noise performance with high reflection gain, we require minimal signal loss via port 2, which demands $\left|S_{23}[\omega_{s}]\right|\ll\left|S_{13}[\omega_{s}]\right|$. In the conventional design with capacitively coupled ports, this condition is satisfied by having a large coupling capacitance at port 1 and a small coupling capacitance at port 2. However, this choice also leads to poor pump transmission from port 2, in conflict with our goal of maximizing pump coupling efficiency.
@@
The frequency separation between the pump and signal for the three-wave-mixing process provides an opportunity to satisfy these seemingly conflicting requirements. In the following sections, we demonstrate that the optimization goals at pump frequency $\omega_{p}$ and signal frequency $\omega_{s}$ can be achieved simultaneously by engineering the coupling network to be strongly frequency selective on a three-wave-mixing JPA.